OpenStudy (anonymous):

(1) If f(x) is a non-constant polynomial satisfying 27f(x^3)−4f(x^2)−x^6 f(3x)+46=0, what is f(2)? (2) a,b and c are positive integers such that f(x)=ax2−bx+c has 2 distinct real roots α and β that satisfy the strict inequality 0<α<β<1. What is the minimum value of a?

OpenStudy (anonymous):

can u PLZ draw this 27f(x3)−4f(x2)−x6f(3x)+46=0

OpenStudy (anonymous):

@saloniiigupta95

OpenStudy (anonymous):

I mean 27f(x3)−4f(x2)−x6f(3x)+46=0 l is this x^3

OpenStudy (anonymous):

Well, I got f(0)=2, f(1)=-1/2 and f(3)=23/2

OpenStudy (anonymous):

f(2) may take a long process to find it.

OpenStudy (anonymous):

but f(0)=-2 not 2